(14 points) College Graduation Rates. Data from the College Results Online website compared the 2011 graduation rate and medi4. Complete the following sentence to interpret the slope of the regression line: in Median SAT score corresponds to a/an inc(14 points) College Graduation Rates. Data from the College Results Online website compared the 2011 graduation rate and median SAT score for 92 similar-sized public universities and colleges in the United States. The scatterplot below shows the relationship between these two variables along with the least squares fit. Round all calculated results to 4 decimal places.
Answer #1

3)

x̅ = 1029.7
y̅ = 48.1
sx= 90.4
sy = 13.9
r= 0.58
slope b1=r*sy/sx= 0.08918
intercept bo=y̅ +r(sy/sx)*(x̅)= -43.730

y^ = -43.7301+0.0892x

4)

an incerase of 1 unit in ,,,,,,,increase of 0.0892 in

5)

predicted value =0.08918*1028-43.7301= 47.9484

6)

residual =48.4-47.9484 = -0.5484

lower than

7)o A. No, because 1455 is beyond the range of the data used to build the regression model.